Nicoley269:

Given: Ray BD bisects angle ABC and measure angle ABC equals 25. Solve for x.

Nicoley269:

1 attachment
Nicoley269:

@dude

dude:

I have an idea on how to do this but I feel there is probably a much more easier way to do this Okay, combine like terms

Nicoley269:

Ok

dude:

What did you get, when combining like terms? xD

Nicoley269:

X-2y=25

dude:

Right, so now isolate x Do you know how to do that?

Nicoley269:

No

dude:

(Do the opposite of subtracting 2y)

Nicoley269:

X=25+2y

dude:

Okay so now we know what x is in terms of y

dude:

Lets go back into our original equation and for every x value, insert 25+2y instead \((2\boxed{x}-y)+(3y-\boxed{x})=25\)

Nicoley269:

Ok....

dude:

\(2(25+2y)-y+3y-(25+2y)=25\) Does that make sense?

Nicoley269:

Yes

dude:

Okay, now distribute the 2 in `2(25+2y)` and -1 in `−(25+2y)` And then combine like terms

Nicoley269:

50+4y-25-2y?

dude:

Right so now you have \(50+4y-y+3y-25-2y=25\) Do you know how to simplify that?

Nicoley269:

4y=0?

dude:

Divide by 4 and you get y=0 Yes thats what I got, so now we substitute 0 into our original equation

dude:

Hmm I made a mistake, let me check my work

dude:

Im a derp ahh

dude:

Screw what I just said

Nicoley269:

All of it?

dude:

Yes, I over-read this

Nicoley269:

Ok...